Week of 170717

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Andrea (MW Officer + data management), Krystof (computing), Maarten (SCOD + ALICE), Paul (storage), Petr (ATLAS), Renato (LHCb), Vincent (security)

  • remote: Alexander (NLT1), David B (IN2P3), David M (FNAL), Di (TRIUMF), Jean-Roch (CMS), John (RAL), Kyle (OSG), Sang-Un (KISTI), Vincenzo (EGI), Xin (BNL)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Activities:
      • stable production with high number of cores provided by HPC
      • trying to optimize number of pilots sent to the CEs (queued & empty pilots)
    • Issues
      • ATLAS CVMFS @ RAL Stratum 1 out of sync for a day
      • BigPanda monitoring web access HTTP -> HTTPS (SSO) - doesn't work for some users

  • CMS reports ( raw view) -
    • Spiky usage of global pool waiting for start of large campaign
    • Issue with site config at RAL GGUS:129562 last Friday, solved Monday morning
    • Transfer issue with JINR GGUS:129528 related to LHCOPN
    • Lots of ghost data removed to close pending transfers, will check within CMS the origin of mis-cataloging.
    • Some minor T2 issues, nothing unusual

  • ALICE -
    • Continued very high activity on average
    • CERN: almost all HTCondor CEs in bad shape during the weekend (GGUS:129571)
      • Tens of thousands of jobs lost
      • Only ~4k job slots available for most of Sunday
      • Fixed Monday morning

  • LHCb reports ( raw view) -
    • Activity
      • Lots of user analysis(some failling) and Monte Carlo simulation
    • Site Issues
      • T0: Jobs hold at HTCondor CEs CERN-PROD (GGUS:129147)
      • T1:
        • NTR

Sites / Services round table:

  • ASGC: NTR
  • BNL: NTR
  • CNAF:
  • EGI: NTR
  • FNAL: NTR
  • IN2P3: NTR
  • JINR:
    • There is some progress in improving the reliability of our link to Moscow, and we continue working on it.
    • We continue working on Ipv6 routing on LHCOPN. Currently, there is a problem with PIC, which has MTU 9000.
  • KISTI: NTR
  • KIT:
  • NDGF: A few pool nodes at UiO needs firmware upgrades. Will be done tomorrow between 13-15 CEST. Some ATLAS data unavailable. (Christian: I will probably have trouble connecting.)
  • NL-T1: Tomorrow SURFsara has an all-day downtime due to network maintenance. We'll be reinstalling the top of rack switches and replacing the qfabric interconnects.
  • NRC-KI:
  • OSG: NTR
  • PIC:
  • RAL: Lots of problems with Castor over the weekend especially for CMS. The castor team are investigating. We need to reboot some disk servers for patching and firmware updates this week.
  • TRIUMF: NTR

  • CERN computing services: HTCondor CE problems during the weekend (OTG:0038736)
  • CERN storage services: NTR
  • CERN databases:
  • GGUS: NTR
  • Monitoring:
    • Availability Report for June 2017 distributed
  • MW Officer: NTR
  • Networks:
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r13 < r12 < r11 < r10 < r9 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r13 - 2017-07-17 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback